OpenStudy (anonymous):

Question 15 Classify the expression 7x3 + 2x2− 9x + 1 as a monomial, binomial, trinomial or polynomial. monomial binomial trinomial polynomial Question 16 Multiply x to the 1 fourth power times x to the 5 eighths power x to the 1 half power x to the 3 eighths power x to the 7 eighths power x to the 5 over 32 power Question 17 Identify all of the following solutions of square root of x plus 8 end root minus 6 equals x. x = −4 x = −7 x = −4 and x= -7 none of the above

OpenStudy (anonymous):

@math&ing001

OpenStudy (math&ing001):

mono = 1st degree bi = 2nd degree tri = 3rd degree

OpenStudy (anonymous):

so its trinomial?

OpenStudy (math&ing001):

Yep !

OpenStudy (anonymous):

#16?

OpenStudy (math&ing001):

\[x ^{a}*x ^{b}=x ^{a+b}\]

OpenStudy (anonymous):

im confuseddddddd

OpenStudy (math&ing001):

Consider a=1/4 and b=5/8

OpenStudy (anonymous):

so , x^5/32

OpenStudy (anonymous):

@math&ing001

OpenStudy (math&ing001):

Do not multiply, add

jimthompson5910 (jim_thompson5910):

7x3 + 2x2− 9x + 1 is not a trinomial. It does NOT have 3 terms. It has more than 3 terms so it is polynomial (poly means many)

OpenStudy (anonymous):

so its x^1/2

OpenStudy (math&ing001):

1/4 + 5/8 = 1/2 ?

OpenStudy (anonymous):

6/12 simplified is 1/2

OpenStudy (math&ing001):

Those are fractions we add them like this : \[\frac{ 1 }{ 4 }+\frac{ 5 }{ 8 }=\frac{ 2+5 }{ 8 }=\frac{ 7 }{ 8 }\]

OpenStudy (anonymous):

where did you get 2+5?

OpenStudy (math&ing001):

And btw I made a mistake be4. As @jim_thompson5910 said monomial, binomial... do not depend on degree but on the number of terms the polynomial has.

OpenStudy (anonymous):

& #17?

OpenStudy (math&ing001):

To add fractions you need to have the same denominator.

OpenStudy (math&ing001):

So 1/4 = 2/8. That's how we got 1/4 + 5/8 = 2/8 + 5/8 = (2+5)/8 = 7/8

OpenStudy (anonymous):

& so it is polynomial ?

OpenStudy (math&ing001):

yeah

OpenStudy (anonymous):

ok & #17?

OpenStudy (math&ing001):

Square root of x plus 8 end root minus 6 equals x. What do you mean by 'end' ?

jimthompson5910 (jim_thompson5910):

"square root of x plus 8 end root minus 6 equals x" looks like it translates to \[\Large \sqrt{x+8} - 6 = x\]

OpenStudy (anonymous):

Yes^

OpenStudy (anonymous):

@math&ing001

OpenStudy (math&ing001):

Get 6 to the other side then put everything to the power of 2 and solve for x. Should look like x+8 = (x+6)^2

OpenStudy (math&ing001):

Oh and check for extraneous solutions !

OpenStudy (math&ing001):

Did you study how to solve quadratics ?

OpenStudy (anonymous):

no.

OpenStudy (math&ing001):

Well you can't do this kind of exercises if you haven't.

OpenStudy (math&ing001):

Last resort is you just plug in the solutions they gave you in choices and see if they check...

OpenStudy (anonymous):

so its -7 ?

OpenStudy (math&ing001):

Check again !

OpenStudy (anonymous):

ok... -4?

OpenStudy (math&ing001):

yep : )

OpenStudy (ranga):

For the first part, a monomial has just one TERM, a binomial has two terms, a trinomial has 3 terms and for more than three terms just polynomial.